OpenStudy (anonymous):

Determine if lines that pass through the given points are parallel, perpendicular or neither. Line A: (5, 4) and (7, 5) Line B: (4, -5) and (8, -3) A. Parallel B. Perpendicular C. Neither D. Cannot be determined

OpenStudy (loser66):

there are many ways to come up, the easiest one is write down the equation of the lines, graph them and look at those graph to give out the answer hope this helps

OpenStudy (anonymous):

\[ \text{ No , the easiest way is To calculate the slope of each lime . then, based on their}\\ \text { relationship, which you should know if you were assigned this problem, determine }\\ \text{whether they have Parallel or Not}\\\]

OpenStudy (anonymous):

\[ m _{1}=\frac{ 5-4}{ 7-5 } =\frac{ 1 }{ 2}\\ m _{2}=\frac{ -3-(-5) }{ 8-4} = \frac{ 1 }{ 2 }\\ \]

OpenStudy (anonymous):

\[\text { hmm , the slopes are equal , I wonder what that means } \]

OpenStudy (loser66):

that means they are skew, isn't it? hehehe